OpenStudy (anonymous):

Byron is 3 years older than Doug. The product of their ages is 40. How old is Doug?

OpenStudy (unklerhaukus):

\[B=D+3\text{ yr}\] \[B\times D=40\text{ yr}\]

OpenStudy (anonymous):

b=3+d; b*d= 40 => (3+d)*d=40 => d^2+3d-40=0 => d= -8, 5. So Doug is 5.

OpenStudy (anonymous):

I still dont get it

OpenStudy (anonymous):

Which part?

OpenStudy (unklerhaukus):

do you understand how to get the two equations from the question/

OpenStudy (anonymous):

b times d is 5 times 8 right

OpenStudy (unklerhaukus):

combining the two equations\[B\times D=40 \text{ yr}\] \[(D+3\text{ yr})\times D=40 \text{ yr}\] \[D^2+3D=40\] \[D^2+3D-40=0\] \[(D+5)(D-8)=0\] \[D=-5,8\]

OpenStudy (anonymous):

still dont get it man smh its hard to understand

OpenStudy (anonymous):

Let Dougs age = x yrs Therefore Byron = (x + 3) yrs Multiply the two ages to get 40: x(x + 3) = 40 x^2 + 3x - 40 = 0 after simplifying (x+8)(x-5) = 0 after factorising x = -8 or x = 5 But age cannot be negative therefore he is % yrs old

OpenStudy (anonymous):

I mean 5 years old (not % yrs old)

OpenStudy (anonymous):

this helps thanks